OpenStudy (hellokitty17):

Please help can't decide which one it is. I have it graphed but I can't tell what it is. WILL MEDAL!!! @phi

OpenStudy (phi):

this looks painful I would first plot the points just to see what it looks like

OpenStudy (phi):

Here is a plot from geogebra

OpenStudy (phi):

by eyeball it looks like it could be a right triangle. To prove it, we would find the slopes from A to B and the slope from B to C can you do that ?

OpenStudy (phi):

find the slope between (1,4) and (3,1) change in y divided by change in x

OpenStudy (phi):

I would do the y value from (1,4) minus the y value from (3,1) can you do that ?

OpenStudy (phi):

now do the x value of (1,4) minus the x value from (3,1)

OpenStudy (phi):

now the slope is 3/-2 or -3/2 next we find the slope from B to C (3,1) and (6,3) can you do that ?

OpenStudy (phi):

first "change in y"

OpenStudy (phi):

and the slope is ⅔ (the minus/minus becomes plus)

OpenStudy (phi):

if the slopes of two lines are "negative reciprocals" that means the lines form a right angle negative reciprocal means if one has slope m, the other has slope -1/m or , (means the same thing) if you multiply the two slopes, you get -1

OpenStudy (phi):

in this case \[ - \frac{3}{2} \cdot \frac{2}{3}= -1 \] proves the lines AB and BC form a 90 degree angle, so the triangle ABC is a right triangle (because it has a 90 degree angle in it)

OpenStudy (phi):

notice I used the graph to guess which lines to check. If we did not have the picture, we would have to check all 3 angles... lots of work